OpenStudy (anonymous):

graph -4x+7y=-28 (show your work)

OpenStudy (shamil98):

Put your equation to slope-intercept form. y = mx + b

OpenStudy (shamil98):

b is the y-intercept from where you start graphing the coefficient m is your slope, rise/run

OpenStudy (anonymous):

so would it be like; 4x+7y=-28 add 4x to both side so 4x+7y=-28+4x then 7y=4x-28 which is y=4/7x-4

OpenStudy (anonymous):

@shamil98